The length of a rectangle is 5 cm less than three times its width. If the perimeter is 70 cm, find the dimensions.
Answer by Maths68(1474)

Width = w=?
Length = l=3w-5
Perimeter = p = 70

Perimeter of a Rectangle = length + length + width + width
Perimeter of a Rectangle =2 length +2 width
P = 2l +2 w
70=2(3w-5)+2w
70=6w-10+2w
70=8w-10
70+10=8w
80=8w
80/8=8w/8
10=w
w=10



Dimensions
============
Width = w=10cm
Length = l=3w-5=3(10)-5=30-5=25cm
Perimeter = p = 70cm